OpenStudy (anonymous):

18 - 64 = 4a - 4a - 4 (a + 3)

OpenStudy (anonymous):

18 - 64 = 4a -4a -4(a +3) -46 = -4a - 12 -34 = -4a a = 34/4 = 17/2

OpenStudy (isaiah.feynman):

 Can you continue from there?